two trucks are travel towards each other from points 1,ooo km apart. The two trucks meet each other in 8 hours. What is the average speed of each truck if one truck travels 15 kilometersper hour faster than other?
S=d/t

Answer by mananth(16946)

Truck I x km/h
Truck II x+ 15
..
They are moving towards each other They are moving towards each other

So add up their speed.

combined speed = x+ 1 x + 15
( 2 x + 15 )
Time = 8 hours
Distance = 1000 miles
Distance = speed * time
( 2 x 15 )* 8 = 1000
16 x + 120 = 1000
16 x = 1000 + -120
16 x = 880
/ 16
x= 55 mph Truck I
Truck II 55 + 15 = 70 mph